OpenStudy (anonymous):

Which equation is equivalent to 2x + 6y = 12? A. y= 1/3x+ 2 B. y= 1/3x-2 c. y= -1/3x+2 d. y= -1/3x -2

OpenStudy (the_fizicx99):

Convert 2x + 6y = 12 to y = mx + b Subtract 2x from both sides Divide every term by 6 to isolate y

OpenStudy (anonymous):

- 1/3 + 2?

OpenStudy (the_fizicx99):

y = -1/3 + 2 yes

OpenStudy (anonymous):

YAY I CAN MATH

OpenStudy (the_fizicx99):

Gratz boi *gives cookie*

OpenStudy (anonymous):

:-) i gave youz a medal. c:

OpenStudy (the_fizicx99):

I see's that